OpenStudy (anonymous):

find the lcm of f^2 , f^2-4 ,and f^2-4f+4

OpenStudy (mathmate):

\(f^2=f*f\) \(f^2-4=(f+2)(f-2)\) \(f^2-4f+4=(f-2)(f-2)\) So you're looking for the simplest polynomial that has all of the above cases as factors.

OpenStudy (anonymous):

@mathmate so would it be f^2 (f+2) (f-2)^2

OpenStudy (mathmate):

Exactly, well-done! But do get familiar with how to factor expressions.

OpenStudy (anonymous):

@mathmate okay, thank you!

OpenStudy (mathmate):

You're welcome! :)